OpenStudy (anonymous):

how do i write 9x+2y=6 in slope inetcept form

OpenStudy (bloomlocke367):

solve for y.. do you know what to do to do that?

OpenStudy (anonymous):

9x + 2y = 6 2y = -9x + 6 y = -9/2x + 3

OpenStudy (anonymous):

thx @creegan and @BloomLocke367

OpenStudy (anonymous):

ur welcome

OpenStudy (bloomlocke367):

you're welcome.